eviant:

PLS HELP!!! Let f(x) = -2x - 7 and g(x) = -4x + 6. Find f*g(-5)

eviant:

@Vocaloid

Vocaloid:

if that's a multiplication sign between f and g, just multiply f(x) and g(x) together then plug in x = 5 into the result

eviant:

yes that is a multiplication sign so, -2(5)-7=-17 -4(5)+6=-14

Vocaloid:

oh, well yeah you can plug in x = 5 first after that you just multiply the results (-17) * (-14)

eviant:

238

Vocaloid:

good so 238 = your solution

eviant:

–6 –59 26 3

eviant:

^^those are the options @Vocaloid

Vocaloid:

oh duh it says x = -5 not x = 5 :S i misread

Vocaloid:

so you'd plug in -2(-5)-7= -4(-5)+6= and multiply the results

eviant:

26*3=78

Vocaloid:

huh. that's not a choice either. hold on... :S

Vocaloid:

I believe it's asking for the function composition not the product so f composed with g, take g(x) and plug it into f(x) to get f(x) = -2x - 7 ---> f(g(x)) = -2(-4x + 6) -7 then plug in x = -5

eviant:

-2(-4(-5)+6)-7=-59